Recent Log Entries

Box Supplies Were Cycled

Log Number: 3814754Logbook(s): Tags: Body: 

The following box supplies were CYCLED by version 2-7 of the Box Supply Cycling Script:
ARC1A Ramp Rate = 8.113 A/s
ARC2A Ramp Rate = 6.895 A/s
ARC3A Ramp Rate = 9.843 A/s
ARC4A Ramp Rate = 10.716 A/s
ARC5A Ramp Rate = 10.166 A/s
ARC6A Ramp Rate = 8.984 A/s
ARC7A Ramp Rate = 10.209 A/s
ARC8A Ramp Rate = 11.956 A/s
ARC9A Ramp Rate = 9.124 A/s
ARCAA Ramp Rate = 12.283 A/s
LAM1C Ramp Rate = 10.484 A/s
LAM3C Ramp Rate = 11.897 A/s
MYA2T01 Ramp Rate N/A
MYA4T01 Ramp Rate N/A
MYA6T01 Ramp Rate N/A
MYAAT01 Ramp Rate N/A
RSEP8A Ramp Rate = 10.471 A/s
RSEP9A Ramp Rate = 9.366 A/s

1L12 cooldown curves

Log Number: 3814753Logbook(s): Entry Makers: beneschBody: 

Top panel: cooldown diodes, #1 is broken.

Middle panel: vacuum readings on ion pumps. 1L1220LOG stuck at 1E-7 Torr is troubling. The other three waveguide pumps are way down.

Bottom panel: per linked entry, LL has been set to 20% max with L tube. Shield is still cooling.

Tomorrow swing or Thursday owl should tell the tale.

FSD to BSY Dump -- Loose Ends

Log Number: 3814749Logbook(s): Entry Makers: moserBody: 

Figure 1: A clear FSD! Figure 2: Lots of caveats to that clearness... So much so that another Figure, Figure 3, needed to capture it all.

First nine anomalies are all related to RF separators. Not powered at the moment, and likely won't be until next week.

Next, BSY dump, IBD4C00 which is a known issue from last run. I verified that indeed, FSD bit for it that sums together a bunch of parameters, Figure 4, still unable to be cleared.

1L07, 1L11, 1L12, 1L13 and the NL C100s are masked at the moment. 1L07 is struggling to stay up, but when it is we verified the FSD for it can be cleared.

1L12 Cooldown is Complete

Log Number: 3814745Logbook(s): Body: 

Ave cavity cooldown rates: 80-100K/hr
All valves have been set to normal operation settings.
LL: 4%
The LL setpoint has been set to 20% due to the L tube installed.
Shield will cooldown over the next day or two. Currently, the shield return temp is 134K.
All pressures check out.
Ins Vac is still being pumped on.

The only thing that I noted during the cooldown is that the midplane diode (R1CXCHET1n) is not reading back correctly. Currently, the diode is reading back 200K, which is incorrect compared to the rest module.

VFV2L02 Reports Open, But Preventing Clear FSD

Log Number: 3814746Logbook(s): Entry Makers: moserBody: 

Shaking the FSD Tree to try and preemptively identify any issues we might have when setting masking for spinning up the main machine.

VFV2L02 appears to be open, along with surround valves. Surrounding pumps are all clear, no faults. Despite this, however, the FSD bit for this valve, summed together with 10 others, FSD_SL03_1, Figure 3, will not clear. Thinking the issue is with VFV2L02 and not the other ten valves because of the red box next to its name, Figure 4. PV for this box = VDISL1180.BE

I don't recall exactly what this box indicates... if I even knew in the first place.

LEM North Linac - failure

Log Number: 3814741Logbook(s): Tags: Body: LEM 9.6 was unsuccessfully run for the North Linac on Tue Jul 14 17:03:07 2020. The operation terminated after 0:08:35 with the following results.
Cavity summarySuccess: 189 Abort: 11 Fail: 0

BSY High Power Dump, ABIL tag #3889, Expected Removal Date changed.

Log Number: 3814737Logbook(s): Body: The following Bypassed Interlock had it's Expected Removal Date changed in the Accelerator Bypassed-Interlocks Log (ABIL) Previous Expected Removal Date: 04 May 2020 New Expected Removal Date: 30 Sep 2020 Who made the change: Daniel Moser Justification for the change: From: DiPette on 2/20/20 -- BSY Dump has been tripping off approx.

Operator SIGN IN

Log Number: 3814728Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Forman
DATE:             07/14/2020
SHIFT:               Swing

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log

FC#2 Unlocked

Log Number: 3814725Logbook(s): Tags: Entry Makers: moserBody: 

ILD still locked in. Forman operator, Moser CC, on-site.

Plan is to finish bunchlength measurements using backphasing method to the 4D spectrometer.

Pages